a + b + c = 180
b = 3a c = 5a a=?

Answer by sharon1948(6)

#287249
a + b + c = 180
b = 3a
c = 5a
a=?
Solve
You know the equation, now plug in the values that b & c equal
a+3a+5a=180
Combine like terms
9a=180
Divide by 9 to isolate a
180/9 = 20
Therefore a = 20